• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 16:17:02 on April 16, 2007

    Entry number 199268

    keyword=Shift Summary DAY -- Monday Apr. 16


    Shift Summary for April 16, 0800-1600


    Shift Crew:

    Shift Leader: Brad Sawatzky
    Target Operator: Lingyan Zhu

    Kinematics:

    Q^2: 55 fm^-2
    Beam Energy: 0.9914 GeV
    Beam Current: 95 uA
    Left HRS: 0.59641 GeV/c @ 140.3 degrees
    Right HRS: 0.72621 GeV/c @ 14.9 degrees

    Summary:


    Time    Description
    0800    Started shift taking production data on He3
    0905    MCC called:  They are going to study Hall A FFB.  They are going
    to use·
            tune beam to start with so we stopped the current run (they're using
            pulsed beam).
    1015    Start taking a little data now that we're back to CW mode (run
    2541).··
            MCC is still working on the FFB though so the data may be junk ( beam
            is trippy and the beam position is pretty erratic).
    1040    Going to Controlled Access (survey started)
    1110    In Controlled Access, Techs working on
            - Left Q1 trip problem
            - change card on Left VDC v2 plane to fix hot wire
    1600    Techs still in the Hall working on Left Q1
    
    

    Run List:


    Run     Target      Events      Comments
    2538    He3         1.4M        Started during last shift
    2539    He3         1.4M        
    2540    He3         0.3M        Stopped early for beam study work
    2541    He3         0.8M        We're back to CW so we'll take a little
    data during the FFB work
                                    Beam position a little more erratic due
    to the FFB work 
    2542    --           --         JUNK (testing) 
    2543    --           --         JUNK (testing)
    2544    --           10k        Pedestal run